Twelve​ samples, each containing five​ parts, were taken from a process that produces steel rods at Emmanual​ Kodzi's factory. The length of each rod in the samples was determined. The results were tabulated and sample means and ranges were computed. The results​ were:

SAMPLE

SAMPLE MEAN

(IN.)

RANGE (IN.) SAMPLE SAMPLE MEAN (IN.) RANGE (IN.)
1 9.104 0.044 7 9.103 0.021
2 9.100 0.051 8 9.103 0.058
3 9.089 0.042 9 9.097 0.039
4 9.108 0.037 10 9.103 0.038
5 9.097 0.048 11 9.103 0.054
6 9.097 0.053 12 9.108 0.061

For the given data, the X (double bar) =______inches (round your response to four decimal places)

Based on the sampling done, the control limits for 3-sigma x-bar chart are:
Upper Control Limit (UCLxbar ) =______ Inches (round your response to four decimal places)
Lower Control Limit (LCLxbar ) =______ Inches (round your response to four decimal places)

Based on the x-bar chart, the process is a) OUT OF CONTROL or b) IN CONTROL

For the given data, the R (average range of the samples) =______ Inches (Round to four decimal places)

The control limits for the 3-sigma R-chart are:
Upper Control Limit (UCLR) =_______inches (round your response to for decimal places)
Lower Control Limit (LCLR) =______ inches (round your response to four decimal places)

Based on the R-chart, the wire cutting process is a) OUT OF CONROL OR b) IN CONTROL
Answer #1

Answer: R bar is the mean of all Range. The formula used is = sum of all range/12.

Similarly, X-bar-bar is mean of all X-bar. The formula used is = sum of all X bar/12.

R= Max- Min X-bar
9.104 0.044
9.1 0.051
9.089 0.042
9.108 0.037
9.097 0.048
9.097 0.053
9.103 0.021
9.103 0.058
9.097 0.039
9.103 0.038
9.103 0.054
9.108 0.061
9.101 0.046
R-bar X-bar-bar

Answers related to X bar

control limits for x-bar
UCL= Xbar-bar+ 3*STD 0.0773
CL= X-bar-bar 0.0455
LCL= Xbarbar- 3*STD 0.0137
standard deviation X-bar= 0.0106
Yes, the process is in control. Because all the points lie within its control limits. And all the points of X-bar lies in X-bar control limits

Answers related to Range

control limits for R
UCL=R-bar+ 3*STD 9.1164
CL= R-bar 9.1010
LCL=R-bar- 3*STD 9.0856
standard deviation R= 0.0051
Yes, the process is in control. Because all the points lie within its control limits. And all the points of range lies in Range control limits

Please ask, if you have any doubts through comment section. Do rate the answer